OpenStudy (anonymous):

Find the slope of the line containing the points (–3, 1) and (3, 4).

myininaya (myininaya):

\[\frac{4-1}{3-(-3)}=\frac{3}{3+3}=\frac{3}{6}=\frac{1}{2}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

rise over runs